Factorise the following expressions
x^2yz+xy^2 z +xyz^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

x^2yz+xy^2 z +xyz^2
x^2yz = x×x×y×z
xy^2z = x×y×y×z
xyz^2= x×y×z×z
Common factor of x^2yz , xy^2z and xyz^2are x, y, z i.e. xyz
Now, x^2yz+xy^2z+xyz^2= xyz(x+y+z)

Thanks